The sum of the present ages of Vatha and Chris is 36. In 4 years time, the sum of their ages
will equal twice Vatha’s present age. How old are they now?

Answer by ankor@dixie-net.com(22740)


v = V's age; c = C's age; write an equation for each statement
:
The sum of the present ages of Vatha and Chris is 36.
v + c = 36
In 4 years time, the sum of their ages will equal twice Vatha’s present age.
(v+4) + (c+4) = 2v
v - 2v + c + 8 = 0
-v + c = -8
Add to the first equation
-v + c = -8
v + c = 36
-----------------Adding eliminates v, find c
0 + 2c = 28
c = 14 is C's age
I'll let you find V's age. check your solutions in both statements
: